zarkam21 (zarkam21):

3. Uranium-238 decays into lead. For a given amount of 238U, half is converted to lead every 4.5 million years. a) If a sample is found with 1/8 (0.125) of the expected 238U, how old would this sample be? (2 points) 13.5 billion b) For carbon-14, 60,000 years is the maximum age that can be accurately dated. What would be the maximum age of a sample if 238U were used? (3 points)

zarkam21 (zarkam21):

@mjdennis

OpenStudy (mjdennis):

OK, what formulas do you have for half-life or decay constant. That's a good place to start.

zarkam21 (zarkam21):

Well I know that half is converted to lead every 4.5 million years.

zarkam21 (zarkam21):

And when I got the answer for part a, I divided.

OpenStudy (mjdennis):

Hmmm. I'm not qualified to help with the physics part. We need to know the half-life of C-14 then. (Half-life of lead is 4.5 million years, because half is converted to lead every 4.5 million years). If you find the half-life of carbon-14, we can get started.

zarkam21 (zarkam21):

6,000 years

zarkam21 (zarkam21):

That is the half-life of carbon-14

OpenStudy (mjdennis):

Got it. So the limit is 60,000/6,000 = 10 half-lives. This is the limit because there is so little of the radioactive stuff left, you can't detect it accurately. Let's find out what fraction of C-14 is left after 10 half-lives: Half is left after 1 half life, so (1/2)^1 = 1/2 is left After 2 half lives, (1/2)^2 is left. Makes sense, right? So let's use the same rule for U-238... How long is 10 half-lives?

zarkam21 (zarkam21):

10

OpenStudy (mjdennis):

how long in billions of years is 10 half-lives of U-238 (read the problem statement again.)

zarkam21 (zarkam21):

1/2^10

OpenStudy (mjdennis):

For 238U, half is left every 4.5 million years. So (1/2)^2 is left in 2*4.5 Myrs, or in 9 million years. (1/2)^3 is left in 3*4.5 Myrs, or in 13.5 million years. (Just like part A) Our ultimate limit was (1/2)^10. How many years for (1/2)^10 ? Just like part A, but 10 times instead of 3 times.

OpenStudy (mjdennis):

I have to go for the day. The answer is on the tip of your tongue. You can get this. This problem, I was probably wrong to have you come over to physics, it was more like math. Sorry about that!

zarkam21 (zarkam21):

45?

OpenStudy (mjdennis):

45 or 45 _million_ ???

zarkam21 (zarkam21):

Well yeah, I know that lol :) But thanks so much for helping me, I always really appreciate it .

OpenStudy (mjdennis):

Glad to help. Just don't take a job as a time-traveler, OK? I'd hate to find out you missed your destination by a million years.

zarkam21 (zarkam21):

What do you mean?

OpenStudy (mjdennis):

It was a joke about our "well, duh" moment over whether you meant "45" or "45 million".

zarkam21 (zarkam21):

Hahaha oh okay :)